The Green Machine Company buys a $250,000 machine to generate revenue of $60,000/yr for five years. At the end of the fifth year the machine is sold for $60,000.What is the net present value of this project? Assume the company wants 10% ROI on investments (their discount rate).

Answer should be in this format in Excel. Functions must be used to solve (IRR, RATE, PV, FV, PMT, etc.)

7) Data Comment yearly investment years of investment (2,500) 20 6% ROI $91,964 Money withdrawn at end
Answer #1

Use EXCEL function PV to calculate the PV of the annual revenues worth $ 60000 and the final sale proceeds worth $ 60000.

In the PV function, rate = 0.1 or 10 % (ROI), nper = 5 years (project duration), FV = $ 60000, PMT = $ 60000 and type = 0 (end of the period)

Using the PV function, we get PV = $ 264702.49

Subtract PV from Initial Machine Cost to arrive at the project's NPV = 264702.49 - 250000 = $ 14702.49